Table 23-6 The table below contains data for the country of Batterland, which produces only waffles and pancakes. The base year is 2013. Price of Waffles $%2.00 $2.00 $2.00 $4.00 Quantity of Waffles 80 100 Price of Pancakes $1.00 $2.00 $3.00 $3.00 Quantity of Pancakes 100 120 150 200 Year 2010 2011 2012 2013 150 13. Refer to Table 23-6. From 2012 to 2013, this countrys output grew a. 28.2%. b, 29.0%. C. 29.6% d. 73.9%

Please explain how you got the percentages

Thank you
Answer #1

Ans: 29.0%

Explanation:

To know the output growth, we have to calculate Real GDP of both the years

Real GDP = (Base year price * Current year quantity)

Real GDP in 2012 = ($4 * 120) + ($3 * 150) = 480 + 450 = $930

Real GDP in 2013 = ($4 * 150) + ($3 * 200) = 600 + 600 = $1200

Country's output growth = [(1200 - 930) / 930] * 100 = 29.0%